OpenStudy (anonymous):

Jordan's Lemma? If anyone here actually knows what it is, help would be appreciated.

OpenStudy (anonymous):

I don't understand it :(

OpenStudy (anonymous):

I just want to know what it is because I need it for the midterm

OpenStudy (anonymous):

@Directrix please help if you can

Directrix (directrix):

Sorry to say: I have not studied it yet.

OpenStudy (anonymous):

Its okay. Thanks for being honest

OpenStudy (michele_laino):

I have a good knowledge about the Jordan's Lemma @Thefaceless

OpenStudy (anonymous):

can you please explain :)

OpenStudy (michele_laino):

please suppose that we have to compute the subsequent integral: \[\int\limits_{ - \infty }^{ + \infty } {f(x){e^{i\nu x}}} dx,\quad \nu \in \mathbb{R}\]

OpenStudy (michele_laino):

and f(x) can be extended to the complex plane, in order to get a holomorphic function f(z)

OpenStudy (michele_laino):

then we can rewrite that integral as follows: \[\int\limits_{ - \infty }^{ + \infty } {f(z){e^{i\nu z}}} dz\]

OpenStudy (anonymous):

ok

OpenStudy (michele_laino):

ok!, now we can demonstrate, that we have the subsequent result: \[\begin{gathered} \int\limits_{ - \infty }^{ + \infty } {f(x){e^{i\nu x}}} dx = 2\pi i\sum\limits_{\operatorname{Im} {z_i} > 0} {{R_f}\left( {{z_i}} \right)} \quad if\;\nu > 0 \hfill \\ \int\limits_{ - \infty }^{ + \infty } {f(x){e^{i\nu x}}} dx = - 2\pi i\sum\limits_{\operatorname{Im} {z_i} < 0} {{R_f}\left( {{z_i}} \right)} \quad if\;\nu < 0 \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

okay i see that

OpenStudy (michele_laino):

where: \[{R_f}\left( {{z_i}} \right)\] is the residual value of the function f(z) at z_i, namely: \[{R_f}\left( {{z_i}} \right) = \frac{1}{{2\pi i}}\oint\limits_\gamma {f\left( z \right)dz} \]

OpenStudy (michele_laino):

being gamma a little circumference in the z-plane which encloses z_i 

OpenStudy (michele_laino):

oops...circumference "gamma"

OpenStudy (anonymous):

okay

OpenStudy (michele_laino):

nevertheless, it is necessary to see how the Jordan's Lemma works, so we can try to compute this integral: \[\int\limits_{ - \infty }^{ + \infty } {\frac{{{e^{ix}}}}{{x\left( {{x^2} + 1} \right)}}} \;dx\]

OpenStudy (anonymous):

wait we haven't shown it yet right?

OpenStudy (anonymous):

this is just preliminary

OpenStudy (michele_laino):

yes!

OpenStudy (michele_laino):

then I'm going to demonstrate the Jordan's Lemma

OpenStudy (anonymous):

okay

OpenStudy (michele_laino):

Let's consider the first case, namely: \[\nu > 0\]

OpenStudy (michele_laino):

then we have to consider an half-circumference which lieas in the upper half-plane, namely: